A rectangular area borders on a stream. A farmer has 400 feet of fencing and intends to use it to fence in the other three sides. Express the area in terms of one of its sides, x.
Answer by edjones(8007)

400/3 = x = one side
x^2 = (400/3)^2 = area
= 160000/9 ft^2
.
Ed